# The complex structure of the Teichmüller space of circle diffeomorphisms in the Zygmund smooth class

### v4

@@ Abstract.
We provide the complex Banach manifold structure for the Teichmüller space of circle diffeomorphisms whose derivatives are in the Zygmund class. This is done by showing that the Schwarzian derivative map is a holomorphic split submersion. We prove similar results for the pre-Schwarzian derivative map and investigate the fiber space over the Teichmüller space. Especially, we show that this projection is also a holomorphic split submersion and induces the structure of a disk bundle. Finally, we introduce the little Zygmund class and consider the closed subspace given by this family of mappings in the Teichmüller space.
@@
Zygmund class, complex Banach manifold structure, holomorphic split submersion, pre-Schwarzian derivative, disk-bundle, little Zygmund
@@ 1. Introduction
The groups of circle diffeomorphisms of certain regularity have been studied in the framework of the theory of the universal Teichmüller space. In the previous papers [10] and [15], we considered the class of circle diffeomorphisms whose derivatives are $\gamma$-Hölder continuous for $10<\gamma<$ 1, and gave the foundation of this Teichmüller space. In the limiting case of this $\gamma$-Hölder continuity as $\gamma\to 1$, one can adopt the Zygmund continuous condition. The corresponding Teichmüller space $T^{Z}$ of those circle diffeomorphisms has been defined in Tang and Wu [14]. In this present paper, we endow $T^{Z}$ with the complex Banach manifold structure and prove several important properties of $T^{Z}$ regarding this complex structure.
@@
${\rm Bel}^{Z}(\mathbb{D}^{*})=\{\mu\in{\rm Bel}(\mathbb{D}^{*})\mid\|\mu\|_{Z}=\underset{\ |z|>1}{\mathrm{ess}\sup}\,((|z|^{2}-1)^{-1}\lor 1)|\mu(z)|<\infty\}.$
It is said that $\mu$ and $\nu$ in ${\rm Bel}(\mathbb{D}^{*})$ are Teichmüller equivalent if the normalized quasiconformal self-homeomorphisms $f^{\mu}$ and $f^{\nu}$ of $\mathbb{D}^{*}$ having $\mu$ and $\nu$ as their complex dilatations coincide on the unit circle $\mathbb{S}=\partial\mathbb{D}^{*}$. The normilization is given by fixing three distinct points on $\mathbb{S}$. The universal Teichmüller space $T$ is the quotient space of ${\rm Bel}(\mathbb{D}^{*})$ by the Teichmüller equivalence. We denote this projection by $\pi:{\rm Bel}(\mathbb{D}^{*})\to T$. Similarly, the Teichmüller space $T^{Z}$ is defined to be the quotient space of ${\rm Bel}^{Z}(\mathbb{D}^{*})$ by the Teichmüller equivalence. This can be identified with the set of all normalized orientation-preserving self-diffeomorphisms $f$ of $\mathbb{S}$ whose derivatives $f^{\prime}$ are continuous and satisfy the following Zygmund condition: there exists some constant $C>0$ such that
@@
for all $\theta\in[0,2\pi)$ and $t>0$. See [14, Theorem 1.1] for these characterizations of $T^{Z}$ as well as those introduced next.
To introduce the complex structure to $T$, we prepare the complex Banach space $A(\mathbb{D})$ of holomorphic functions $\varphi$ on the unit disk $\mathbb{D}=\{z\mid|z|<1\}$ with
$\|\varphi\|_{A}=\sup_{|z|<1}\,(1-|z|^{2})^{2}|\varphi(z)|<\infty.$
In the case of $T^{Z}$, we consider the complex Banach space
$A^{Z}(\mathbb{D})=\{\varphi\in A(\mathbb{D})\mid\|\varphi\|_{A^{Z}}=\sup_{|z|<1}\,(1-|z|^{2})|\varphi(z)|<\infty\}$
as the corresponding space. It is obvious that $\|\varphi\|_{A}\leq\|\varphi\|_{A^{Z}}$.
For every $\mu\in{\rm Bel}(\mathbb{D}^{*})$, let $f_{\mu}$ be the quasiconformal homeomorphism of the extended complex plane $\widehat{\mathbb{C}}$ with complex dilatation 00 on $\mathbb{D}$ and $\mu$ on $\mathbb{D}^{*}$. Then, the Schwarzian derivative map $\Phi$ is defined by the correspondence of $\mu$ to $S_{f_{\mu}|_{\mathbb{D}}}$, where for a locally univalent meromorphic function $f:\mathbb{D}\to\widehat{\mathbb{C}}$, the Schwarzian derivative of $f$ is defined by
@@
It is known that $S_{f_{\mu}|_{\mathbb{D}}}$ belongs to $A(\mathbb{D})$ and $\Phi:{\rm Bel}(\mathbb{D}^{*})\to A(\mathbb{D})$ is a holomorphic split submersion (see [12, Section 3.4]). Moreover, $\Phi$ projects down to a well-defined injection $\alpha:T\to A(\mathbb{D})$ such that $\alpha\circ\pi=\Phi$. This is called the Bers embedding of $T$. By the property of $\Phi$, we see that $\alpha$ is a homeomorphism onto the image. This provides $T$ with the complex Banach structure of $A(\mathbb{D})$, which is the unique complex structure on $T$ such that the Teichmüller projection $\pi:{\rm Bel}(\mathbb{D}^{*})\to T$ is a holomorphic map with surjective derivatives at all points of ${\rm Bel}(\mathbb{D}^{*})$.
For the Teichmüller space $T^{Z}$, it was proved in [14, Theorem 2.8] that
$\Phi({\rm Bel}^{Z}(\mathbb{D}^{*}))=A^{Z}(\mathbb{D})\cap\Phi({\rm Bel}(\mathbb{D}^{*})).$
@@
Theorem 1.1 ([14, Theorem 1.2]).
The Schwarzian derivative map $\Phi:{\rm Bel}^{Z}(\mathbb{D}^{*})\to A^{Z}(\mathbb{D})$ is holomorphic. Hence, the Bers embedding $\alpha:T^{Z}\to A^{Z}(\mathbb{D})$ is continuous.
Indeed, as the Teichmüller projection $\pi:{\rm Bel}^{Z}(\mathbb{D}^{*})\to T^{Z}$ is a quotient map satisfying $\alpha\circ\pi=\Phi$, the continuity of $\alpha$ is equivalent to that of $\Phi$.
In this paper, as the universal Teichmüller space $T$ and certain other spaces are in the case, we improve these results.
Theorem 1.2.
The following are hold true:
@@
    The Schwarzian derivative map $\Phi:{\rm Bel}^{Z}(\mathbb{D}^{*})\to A^{Z}(\mathbb{D})$ is a holomorphic split submersion onto the image $\Phi({\rm Bel}^{Z}(\mathbb{D}^{*}))$.
@@
    The Bers embedding $\alpha:T^{Z}\to A^{Z}(\mathbb{D})$ is a homeomorphism onto the image $\alpha(T^{Z})=\Phi({\rm Bel}^{Z}(\mathbb{D}^{*}))$, which is a connected open subset of $A^{Z}(\mathbb{D})$.
@@
    The Teichmüller space $T^{Z}$ is endowed with the unique complex Banach manifold structure such that the Teichmüller projection $\pi:{\rm Bel}^{Z}(\mathbb{D}^{*})\to T^{Z}$ is a holomorphic map with surjective derivatives at all points of ${\rm Bel}^{Z}(\mathbb{D}^{*})$.
These claims are shown in Corollaries 2.5, 2.6, and 2.7, respectively, as the consequences from Theorem 2.4 in Section 2, which supplies the property of split submersion to the holomorphic map $\Phi$ given in Theorem 1.1.
Next, we consider a certain fiber space $\widetilde{\mathcal{T}}^{Z}$ over the Teichmüller space $\alpha(T^{Z})\cong T^{Z}$, which is realized in the space of pre-Schwarzian derivatives of $f_{\mu}|_{\mathbb{D}}$. Here, for a locally univalent holomorphic function $f:\mathbb{D}\to\mathbb{C}$, the pre-Schwarzian derivative of $f$ is defined by $P_{f}=(\log f^{\prime})^{\prime}$ which satisfies $S_{f}=(P_{f})^{\prime}-(P_{f})^{2}/2$. In Section 3, we are concerned with this fiber space $\widetilde{\mathcal{T}}^{Z}$ and the structure of the projection $\Lambda:\widetilde{\mathcal{T}}^{Z}\to\alpha(T^{Z})$.
For $\mu\in{\rm Bel}(\mathbb{D}^{*})$, we normalize $f_{\mu}$ so that it fixes $\infty$. Then, $\psi=\log(f_{\mu}|_{\mathbb{D}})^{\prime}$ belongs to the space $B(\mathbb{D})$ of all Bloch functions on $\mathbb{D}$, and the pre-Schwarzian derivative map $\Psi:{\rm Bel}(\mathbb{D}^{*})\to B(\mathbb{D})$ is defined by $\mu\mapsto\log(f_{\mu}|_{\mathbb{D}})^{\prime}$, which is known to be holomorphic. The corresponding space to ${\rm Bel}^{Z}(\mathbb{D}^{*})$ is given as
$B^{Z}(\mathbb{D})=\{\psi\in B(\mathbb{D})\mid\|\psi\|_{B^{Z}}=|\psi^{\prime}(0)|+\sup_{|z|<1}\,(1-|z|^{2})|\psi^{\prime\prime}(z)|<\infty\}.$
It was proved in [14, Theorem 1.3] that the pre-Schwarzian derivative map $\Psi:{\rm Bel}^{Z}(\mathbb{D}^{*})\to B^{Z}(\mathbb{D})$ is holomorphic. The fiber space $\widetilde{\mathcal{T}}^{Z}$ is defined as the image of $\Psi$ in $B^{Z}(\mathbb{D})$, and the projection $\Lambda$ onto $\alpha(T^{Z})\subset A^{Z}(\mathbb{D})$ is given by $\Lambda(\psi)=\psi^{\prime}-\psi^{2}/2$. Regarding the projection $\Lambda$, what we obtain in Section 3 are summarized as follows.
Theorem 1.3.
$\Lambda:\widetilde{\mathcal{T}}^{Z}\to\alpha(T^{Z})$ is a holomorphic split submersion, and $\widetilde{\mathcal{T}}^{Z}$ is a topological disk bundle over the Teichmüller space $T^{Z}$ with the projection $\Lambda$.
These results are proved in Theorems 3.3 and 3.4 separately. The first one is a consequence of Theorem 1.2. The second one is a basic claim on this fiber space but it has not been considered properly in the literature even for other Teichmüller spaces.
Finally in Section 4, we define a closed subspace of $T^{Z}$ in a canonical way. The universal Teichmüller space $T$ includes the little subspace $T_{0}$ consisting of all asymptotically conformal classes, which means that $T_{0}$ is defined by the Beltrami coefficients $\mu$ tending to 00 near the boundary. Similarly, the little subspace $T^{Z}_{0}$ of $T^{Z}$ can be defined by the Beltrami coefficients in ${\rm Bel}^{Z}_{0}(\mathbb{D}^{*})$, which consists of $\mu\in{\rm Bel}^{Z}(\mathbb{D}^{*})$ with the condition that $(|z|^{2}-1)^{-1}\mu(z)$ tends to 00 near the boundary. Namely, $T^{Z}_{0}=\pi({\rm Bel}^{Z}_{0}(\mathbb{D}^{*}))$. This corresponds to the subspace consisting of all normalized self-diffeomorphisms $f$ of $\mathbb{S}$ whose continuous derivative $f^{\prime}$ satisfies the Zygmund condition in addition with
$|f^{\prime}(e^{i(\theta+t)})-2f^{\prime}(e^{i\theta})+f^{\prime}(e^{i(\theta-t)})|=o(t)$
uniformly as $t\to 0$. In Section 4, we state some of our results also for the little subspace $T^{Z}_{0}$ and $\alpha(T^{Z}_{0})=\Phi({\rm Bel}^{Z}_{0}(\mathbb{D}^{*}))$, and for the fiber space $\widetilde{\mathcal{T}}^{Z}_{0}=\Psi({\rm Bel}^{Z}_{0}(\mathbb{D}^{*}))$.
We do not deal with all results mentioned in Sections 2 and 3 in this modification; only basic claims are proved in Theorems 4.1 and 4.3. However, because the arguments are almost the same, we can easily obtain other results as well. For instance, the version of Theorem 1.3 can be read as follows and we mention it here without proof.
Theorem 1.4.
$\Lambda:\widetilde{\mathcal{T}}^{Z}_{0}\to\alpha(T^{Z}_{0})$ is a holomorphic split submersion, and $\widetilde{\mathcal{T}}^{Z}_{0}$ is a topological disk bundle over the Teichmüller space $T^{Z}_{0}$ with the projection $\Lambda$.
@@ 2. The complex Banach structure
In this section, we prove Theorem 1.2. This will be done by the combination of Theorem 1.1 and Theorem 2.4, which is the main achievement in this paper. In the first half of this section, we show necessary claims towards this goal.
@@
Proposition 2.1.
If $\mu,\nu\in{\rm Bel}^{Z}(\mathbb{D}^{*})$, then $\mu\ast\nu^{-1}\in{\rm Bel}^{Z}(\mathbb{D}^{*})$. Moreover, for every $\nu\in{\rm Bel}^{Z}(\mathbb{D}^{*})$, the right translation $r_{\nu}:{\rm Bel}^{Z}(\mathbb{D}^{*})\to{\rm Bel}^{Z}(\mathbb{D}^{*})$ defined by $\mu\mapsto\mu\ast\nu^{-1}$ is continuous.
@@
for a constant $C>0$ depending only on $\|\nu\|_{\infty}$. This implies that $(\mu\ast\nu^{-1})\circ f^{\nu}\in{\rm Bel}^{Z}(\mathbb{D})$. By [10, Theorem 6.4], we see that $1-|f^{\nu}(z)|^{2}\asymp 1-|z|^{2}$ (the symbol $\asymp$ stands for the equality modulo a uniform positive constant multiple), where the multiple constant depends only on $\nu$. Hence,
@@
for $w=f^{\nu}(z)$. This shows that $\mu\ast\nu^{-1}\in{\rm Bel}^{Z}(\mathbb{D})$.
Similarly, for any $\mu_{1},\mu_{2}\in{\rm Bel}^{Z}(\mathbb{D}^{*})$, we have
@@
This shows that the right translation $r_{\nu}$ is continuous for every $\nu\in{\rm Bel}^{Z}(\mathbb{D}^{*})$. ∎
@@
Lemma 2.2.
For every $\nu\in{\rm Bel}^{Z}(\mathbb{D}^{*})$, the right translation $r_{\nu}$ is a biholomorphic automorphism of ${\rm Bel}^{Z}(\mathbb{D}^{*})$.
@@
This is a standard consequence from Proposition 2.1. We know that $r_{\nu}:{\rm Bel}(\mathbb{D}^{*})\to{\rm Bel}(\mathbb{D}^{*})$ is holomorphic in the norm of ${\rm Bel}(\mathbb{D}^{*})$. Then, the continuity of $r_{\nu}$ in the norm of ${\rm Bel}^{Z}(\mathbb{D}^{*})$ implies the holomorphy. See [8, p.206]. Proposition 2.1 also implies that $r_{\nu^{-1}}$ is continuous. As $r_{\nu^{-1}}=(r_{\nu})^{-1}$, we see that $(r_{\nu})^{-1}$ is also holomorphic, and that $r_{\nu}$ is biholomorphic. ∎
@@
The right translation $r_{\nu}$ for every $\nu\in{\rm Bel}^{Z}(\mathbb{D}^{*})$ is projected down under $\pi:{\rm Bel}^{Z}(\mathbb{D}^{*})\to T^{Z}$ to a well-defined bijection $R_{[\nu]}:T^{Z}\to T^{Z}$ which depends only on the Teichmüller equivalence class $[\nu]\in T^{Z}$. After obtaining Theorem 2.4 and Corollary 2.7 below, we see that $R_{[\nu]}$ is a biholomorphic automorphism of $T^{Z}$.
@@
Lemma 2.3.
For every $\mu\in{\rm Bel}^{Z}(\mathbb{D}^{*})$, there exists $\nu\in{\rm Bel}^{Z}(\mathbb{D}^{*})$ such that $\nu$ is Teichmüller equivalent to $\mu$ and $f^{\nu}$ is a bi-Lipschitz real-analytic diffeomorphism in the hyperbolic metric on $\mathbb{D}^{*}$.
@@
Suppose that $\|\mu\|_{\infty}<1/3$. Then, $\varphi=\Phi(\mu)$ satisfies $\|\varphi\|_{A}<2$ (see [8, Section II.3.3]). The Ahlfors–Weill section $\sigma$ for the Schwarzian derivative map $\Phi$ is defined by
@@
for $z^{*}=1/\bar{z}\in\mathbb{D}^{*}$ (see [8, Section II.5.1]). Let $\nu=\sigma(\varphi)$. Then, $\nu$ is Teichmüller equivalent to $\mu$ due to $\Phi\circ\sigma={\rm id}$. Moreover, $\nu$ belongs to ${\rm Bel}^{Z}(\mathbb{D}^{*})$ because $\varphi\in A^{Z}(\mathbb{D})$ by Theorem 1.1, and $f^{\nu}$ is a bi-Lipschitz real-analytic self-diffeomorphism of $\mathbb{D}^{*}$ in the hyperbolic metric by [11, Theorem 8]. A similar claim can be found in [13, p.27].
For an arbitrary $\mu\in{\rm Bel}(\mathbb{D}^{*})$, by choosing $n\in\mathbb{N}$ so that $n\geq 3\|\mu\|_{\infty}/(1-\|\mu\|_{\infty})$, we set $\mu_{k}=k\mu/n$ $(k=0,1,\ldots,n)$. Then, $\|\mu_{k+1}\ast\mu_{k}^{-1}\|_{\infty}<1/3$ is satisfied. We will prove the desired claim by induction. Suppose that we have obtained $\nu_{k}\in{\rm Bel}^{Z}(\mathbb{D}^{*})$ such that $\nu_{k}$ is Teichmüller equivalent to $\mu_{k}$ and $f^{\nu_{k}}$ is bi-Lipschitz. By Proposition 2.1, $\mu_{k+1}\ast\nu_{k}^{-1}$ is in ${\rm Bel}^{Z}(\mathbb{D}^{*})$, which is Teichmüller equivalent to $\mu_{k+1}\ast\mu_{k}^{-1}$. Hence,
$\Phi(\mu_{k+1}\ast\nu_{k}^{-1})=\Phi(\mu_{k+1}\ast\mu_{k}^{-1})\in A^{Z}(\mathbb{D}),$
and the argument in the first part of this proof implies that $\sigma(\Phi(\mu_{k+1}\ast\nu_{k}^{-1}))\in{\rm Bel}^{Z}(\mathbb{D}^{*})$ and this yields a bi-Lipschitz self-diffeomorphism of $\mathbb{D}^{*}$. Then, its composition with $f^{\nu_{k}}$ is also a bi-Lipschitz diffeomorphism whose complex dilatation is defined to be $\nu_{k+1}$. Namely,
@@
This is Teichmüller equivalent to $\mu_{k+1}$ and belongs to ${\rm Bel}^{Z}(\mathbb{D}^{*})$ by Proposition 2.1. Thus, the induction step proceeds. We obtain $\nu=\nu_{n}$ as a required replacement of $\mu$. ∎
@@
For circle diffeomorphisms whose derivatives are $\gamma$-Hölder continuous for $10<\gamma<$ 1, the conformally barycentric extension (the Douady–Earle extension) yields an appropriate bi-Lipschitz diffeomorphism (see [10, Theorem 6.10]), but we do not know whether this works also for the Zygmund smooth class. The Beurling–Ahlfors extension of self-diffeomorphisms of $\mathbb{R}$ whose derivatives are in the Zygmund smooth class is investigated in [7].
@@
Theorem 2.4.
For every $\mu\in{\rm Bel}^{Z}(\mathbb{D}^{*})$, there exists a holomorphic map $s_{\mu}:V_{\varphi}\to{\rm Bel}^{Z}(\mathbb{D}^{*})$ defined on some neighborhood $V_{\varphi}$ of $\varphi=\Phi(\mu)$ in $A^{Z}(\mathbb{D})$ such that $s_{\mu}(\varphi)=\mu$ and $\Phi\circ s_{\mu}$ is the identity on $V_{\varphi}$.

### v5

@@ Abstract.
We provide the complex Banach manifold structure for the Teichmüller space of circle diffeomorphisms whose derivatives are in the Zygmund class. This is done by showing that the Schwarzian derivative map is a holomorphic split submersion.
@@
Zygmund class, complex Banach manifold structure, holomorphic split submersion, pre-Schwarzian derivative
@@ 1. Introduction
The groups of circle diffeomorphisms of certain regularity have been studied in the framework of the theory of the universal Teichmüller space. In the previous papers [6] and [11], we considered the class of circle diffeomorphisms whose derivatives are $\gamma$-Hölder continuous for $10<\gamma<$ 1, and gave the foundation of this Teichmüller space. In the limiting case of this $\gamma$-Hölder continuity as $\gamma\to 1$, one can adopt the Zygmund continuous condition. The corresponding Teichmüller space $T_{Z}$ of those circle diffeomorphisms has been defined in Tang and Wu [10]. In this present paper, we endow $T_{Z}$ with the complex Banach manifold structure and prove several important properties of $T_{Z}$ regarding this complex structure.
@@
${\rm Bel}_{Z}(\mathbb{D}^{*})=\{\mu\in{\rm Bel}(\mathbb{D}^{*})\mid\|\mu\|_{Z}={\rm ess\,sup}_{|z|>1}((|z|^{2}-1)^{-1}\lor 1)|\mu(z)|<\infty\}.$
It is said that $\mu$ and $\nu$ in ${\rm Bel}(\mathbb{D}^{*})$ are Teichmüller equivalent if the normalized quasiconformal self-homeomorphisms $f^{\mu}$ and $f^{\nu}$ of $\mathbb{D}^{*}$ having $\mu$ and $\nu$ as their complex dilatations coincide on the unit circle $\mathbb{S}=\partial\mathbb{D}^{*}$. The universal Teichmüller space $T$ is the quotient space of ${\rm Bel}(\mathbb{D}^{*})$ by the Teichmüller equivalence. We denote this projection by $\pi:{\rm Bel}(\mathbb{D}^{*})\to T$. Similarly, the Teichmüller space $T_{Z}$ is defined to be the quotient space of ${\rm Bel}_{Z}(\mathbb{D}^{*})$ by the Teichmüller equivalence. This can be identified with the set of all normalized orientation-preserving self-diffeomorphisms $f$ of $\mathbb{S}$ whose derivatives $f^{\prime}$ are continuous and satisfy the following Zygmund condition: there exists some constant $C>0$ such that
@@
for all $\theta\in[0,2\pi)$ and $t>0$. See [10, Theorem 1.1] for these characterizations of $T_{Z}$.
To introduce the complex structures to $T$, we prepare the complex Banach space $A(\mathbb{D})$ of holomorphic functions $\varphi$ on the unit disk $\mathbb{D}=\{z\mid|z|<1\}$ with
$\|\varphi\|_{A}={\rm sup}_{|z|<1}(1-|z|^{2})^{2}|\varphi(z)|<\infty.$
In the case of $T_{Z}$, we consider the complex Banach space
$A_{Z}(\mathbb{D})=\{\varphi\in A(\mathbb{D})\mid\|\varphi\|_{A_{Z}}={\rm sup}_{|z|<1}(1-|z|^{2})|\varphi(z)|<\infty\}$
as the corresponding space. It is obvious that $\|\varphi\|_{A}\leq\|\varphi\|_{A_{Z}}$.
For every $\mu\in{\rm Bel}(\mathbb{D}^{*})$, let $f_{\mu}$ be the quasiconformal homeomorphism of the extended complex plane $\widehat{\mathbb{C}}$ with complex dilatation 00 on $\mathbb{D}$ and $\mu$ on $\mathbb{D}^{*}$. Then, the Schwarzian derivative map $\Phi$ is defined by the correspondence of $\mu$ to $S_{f_{\mu}|_{\mathbb{D}}}$, where for a locally univalent holomorphic function $f$ in general, the Schwarzian derivative of $f$ is defined by
@@
It is known that $S_{f_{\mu}|_{\mathbb{D}}}$ belongs to $A(\mathbb{D})$ and $\Phi:{\rm Bel}(\mathbb{D}^{*})\to A(\mathbb{D})$ is a holomorphic split submersion (see [8, Section 3.4]). Moreover, $\Phi$ projects down to a well-defined injection $\alpha:T\to A(\mathbb{D})$ such that $\alpha\circ\pi=\Phi$. This is called the Bers embedding of $T$. By the property of $\Phi$, we see that $\alpha$ is a homeomorphism onto the image. This provides $T$ with the complex Banach structure of $A(\mathbb{D})$, which is the unique complex structure on $T$ such that the Teichmüller projection $\pi:{\rm Bel}(\mathbb{D}^{*})\to T$ is a holomorphic map with surjective derivatives at all points of ${\rm Bel}(\mathbb{D}^{*})$.
For the Teichmüller space $T_{Z}$, it was proved in [10, Theorem 2.8] that
$\Phi({\rm Bel}_{Z}(\mathbb{D}^{*}))=A_{Z}(\mathbb{D})\cap\Phi({\rm Bel}(\mathbb{D}^{*})).$
@@
Theorem 1 ([10, Theorem 1.2]).
The Schwarzian derivative map $\Phi:{\rm Bel}_{Z}(\mathbb{D}^{*})\to A_{Z}(\mathbb{D})$ is holomorphic.
Corollary 2.
The Bers embedding $\alpha:T_{Z}\to A_{Z}(\mathbb{D})$ is continuous.
Indeed, as the Teichmüller projection $\pi:{\rm Bel}_{Z}(\mathbb{D}^{*})\to T_{Z}$ is a quotient map satisfying $\alpha\circ\pi=\Phi$, the continuity of $\alpha$ is equivalent to that of $\Phi$.
In this paper, these results are improved as is the case for T and many other Teichmüller spaces.
Theorem 3.
The following hold true:
@@
    The Schwarzian derivative map $\Phi:{\rm Bel}_{Z}(\mathbb{D}^{*})\to A_{Z}(\mathbb{D})$ is a holomorphic split submersion onto the image.
@@
    The Bers embedding $\alpha:T_{Z}\to A_{Z}(\mathbb{D})$ is a homeomorphism onto the image $\alpha(T_{Z})=\Phi({\rm Bel}_{Z}(\mathbb{D}^{*}))$, which is a connected open subset of $A_{Z}(\mathbb{D})$.
@@
    The Teichmüller space $T_{Z}$ is endowed with the unique complex Banach manifold structure such that the Teichmüller projection $\pi:{\rm Bel}_{Z}(\mathbb{D}^{*})\to T_{Z}$ is a holomorphic map with surjective derivatives at all points of ${\rm Bel}_{Z}(\mathbb{D}^{*})$.
These claims are shown in Corollaries 8, 9, and 10, respectively, as the consequences from Theorem 7 in Section 2, which supplies the property of split submersion to the holomorphic map $\Phi$ given in Theorem 1.
The Teichmüller space $T_{Z}$ can be also embedded into the space of pre-Schwarzian derivatives of $f_{\mu}|_{\mathbb{D}}$ instead of using Schwarzian derivatives. This was also investigated in [10] and it was proved that this correspondence to $\mu$ is holomorphic. In Section 3, we also improve the results for this model of $T_{Z}$.
@@ 2. The complex Banach structure
In this section, we prove Theorem 3. This will be done by the combination of Theorem 1 and Theorem 7, which is the main achievement in this paper. In the first half of this section, we show necessary claims towards this goal.
@@
Proposition 4.
If $\mu,\nu\in{\rm Bel}_{Z}(\mathbb{D}^{*})$, then $\mu\ast\nu^{-1}\in{\rm Bel}_{Z}(\mathbb{D}^{*})$. Moreover, for every $\nu\in{\rm Bel}_{Z}(\mathbb{D}^{*})$, the right translation $r_{\nu}:{\rm Bel}_{Z}(\mathbb{D}^{*})\to{\rm Bel}_{Z}(\mathbb{D}^{*})$ defined by $\mu\mapsto\mu\ast\nu^{-1}$ is continuous.
@@
for a constant $C>0$ depending only on $\|\nu\|_{\infty}$. This implies that $(\mu\ast\nu^{-1})\circ f^{\nu}\in{\rm Bel}_{Z}(\mathbb{D})$. By [6, Theorem 6.4], we see that $1-|f^{\nu}(z)|^{2}\asymp 1-|z|^{2}$ (the symbol $\asymp$ stands for the equality modulo a uniform positive constant multiple), where the multiple constant depends only on $\nu$. Hence,
@@
for $w=f^{\nu}(z)$. This shows that $\mu\ast\nu^{-1}\in{\rm Bel}_{Z}(\mathbb{D})$.
Similarly, for any $\mu_{1},\mu_{2}\in{\rm Bel}_{Z}(\mathbb{D}^{*})$, we have
@@
This shows that the right translation $r_{\nu}$ is continuous for every $\nu\in{\rm Bel}_{Z}(\mathbb{D}^{*})$. ∎
@@
Lemma 5.
For every $\nu\in{\rm Bel}_{Z}(\mathbb{D}^{*})$, the right translation $r_{\nu}$ is a biholomorphic automorphism of ${\rm Bel}_{Z}(\mathbb{D}^{*})$.
@@
This is a standard consequence from Proposition 4. We know that $r_{\nu}:{\rm Bel}(\mathbb{D}^{*})\to{\rm Bel}(\mathbb{D}^{*})$ is holomorphic in the norm of ${\rm Bel}(\mathbb{D}^{*})$. Then, the continuity of $r_{\nu}$ in the norm of ${\rm Bel}_{Z}(\mathbb{D}^{*})$ implies the holomorphy. See [5, p.206]. Proposition 4 also implies that $r_{\nu^{-1}}$ is continuous. As $r_{\nu^{-1}}=(r_{\nu})^{-1}$, we see that $(r_{\nu})^{-1}$ is also holomorphic, and that $r_{\nu}$ is biholomorphic. ∎
@@
The right translation $r_{\nu}$ for every $\nu\in{\rm Bel}_{Z}(\mathbb{D}^{*})$ is projected down under $\pi:{\rm Bel}_{Z}(\mathbb{D}^{*})\to T_{Z}$ to a well-defined bijection $R_{[\nu]}:T_{Z}\to T_{Z}$ which depends only on the Teichmüller equivalence class $[\nu]\in T_{Z}$. After obtaining Theorem 7 and Corollary 10 below, we see that $R_{[\nu]}$ is a biholomorphic automorphism of $T_{Z}$.
@@
Lemma 6.
For every $\mu\in{\rm Bel}_{Z}(\mathbb{D}^{*})$, there exists $\nu\in{\rm Bel}_{Z}(\mathbb{D}^{*})$ such that $\nu$ is Teichmüller equivalent to $\mu$ and $f^{\nu}$ is a bi-Lipschitz real-analytic diffeomorphism in the hyperbolic metric on $\mathbb{D}^{*}$.
@@
Suppose that $\|\mu\|_{A}<1/3$. Then, $\varphi=\Phi(\mu)$ satisfies $\|\varphi\|_{A}<2$ (see [5, Section II.3.3]). The Ahlfors–Weill section $\sigma$ for the Schwarzian derivative map $\Phi$ is defined by
@@
for $z^{*}=1/\bar{z}\in\mathbb{D}^{*}$ (see [5, Section II.5.1]). Let $\nu=\sigma(\varphi)$. Then, $\nu$ is Teichmüller equivalent to $\mu$ due to $\Phi\circ\sigma={\rm id}$. Moreover, $\nu$ belongs to ${\rm Bel}_{Z}(\mathbb{D}^{*})$ because $\varphi\in A_{Z}(\mathbb{D})$ by Theorem 1, and $f^{\nu}$ is a bi-Lipschitz real-analytic self-diffeomorphism of $\mathbb{D}^{*}$ in the hyperbolic metric by [7, Theorem 8]. A similar claim can be found in [9, p.27].
For an arbitrary $\mu\in{\rm Bel}(\mathbb{D}^{*})$, by choosing $n\in\mathbb{N}$ so that $n\geq 3\|\mu\|_{\infty}/(1-\|\mu\|_{\infty})$, we set $\mu_{k}=k\mu/n$ $(k=0,1,\ldots,n)$. Then, $\|\mu_{k+1}\ast\mu_{k}^{-1}\|_{\infty}<1/3$ is satisfied. We will prove the desired claim by induction. Suppose that we have obtained $\nu_{k}\in{\rm Bel}_{Z}(\mathbb{D}^{*})$ such that $\nu_{k}$ is Teichmüller equivalent to $\mu_{k}$ and $f^{\nu_{k}}$ is bi-Lipschitz. By Proposition 4, $\mu_{k+1}\ast\nu_{k}^{-1}$ is in ${\rm Bel}_{Z}(\mathbb{D}^{*})$, which is Teichmüller equivalent to $\mu_{k+1}\ast\mu_{k}^{-1}$. Hence,
$\Phi(\mu_{k+1}\ast\nu_{k}^{-1})=\Phi(\mu_{k+1}\ast\mu_{k}^{-1})\in A_{Z}(\mathbb{D}),$
and the argument in the first part of this proof implies that $\sigma(\Phi(\mu_{k+1}\ast\nu_{k}^{-1}))\in{\rm Bel}_{Z}(\mathbb{D}^{*})$ and this yields a bi-Lipschitz self-diffeomorphism of $\mathbb{D}^{*}$. Then, its composition with $f^{\nu_{k}}$ is also a bi-Lipschitz diffeomorphism whose complex dilatation is defined to be $\nu_{k+1}$. Namely,
@@
This is Teichmüller equivalent to $\mu_{k+1}$ and belongs to ${\rm Bel}_{Z}(\mathbb{D}^{*})$ by Proposition 4. Thus, the induction step proceeds. We obtain $\nu=\nu_{n}$ as a required replacement of $\mu$. ∎
@@
For circle diffeomorphisms whose derivatives are $\gamma$-Hölder continuous for $10<\gamma<$ 1, the conformally barycentric extension (the Douady–Earle extension) yields an appropriate bi-Lipschitz diffeomorphism (see [6, Theorem 6.10]), but we do not know whether this works also for the Zygmund smooth class. The Beurling–Ahlfors extension of self-diffeomorphisms of R whose derivatives are in the Zygmund smooth class is investigated in [4].
@@
Theorem 7.
For every $\mu\in{\rm Bel}_{Z}(\mathbb{D}^{*})$, there exists a holomorphic map $s_{\mu}:V_{\varphi}\to{\rm Bel}_{Z}(\mathbb{D}^{*})$ defined on some neighborhood $V_{\varphi}$ of $\varphi=\Phi(\mu)$ in $A_{Z}(\mathbb{D})$ such that $s_{\mu}(\varphi)=\mu$ and $\Phi\circ s_{\mu}$ is the identity on $V_{\varphi}$.